OpenStudy (anonymous):

x/2+ x/3= 5 I know x= 6 but can you show the steps needed I am confused on this one.

jimthompson5910 (jim_thompson5910):

Multiply EVERY term by the LCD 6 to go from x/2 + x/3 = 5 to 3x + 2x = 30 Now solve for x 3x + 2x = 30 5x = 30 x = 30/5 x = 6

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome